OpenStudy (anonymous):

Small pebble thrown horizantally at 20ms/1 from top of cliff 80m high. Ignoring air Res. , and taking g=10ms/2 , Cal: a) how long pebble takes to reach ground b) distance from foot of the cliff to where pebble hits ground. c) Vertical velocity of pebble as it hits ground d) horizontal velocity of pebble as it hits ground.

OpenStudy (anonymous):

a) v = sqrt(2gh) = sqrt(2 x 10 x 80) = sqrt(1600) = 40 mls v = gt so t = v/g = 40/10 t = 4 seconds

OpenStudy (anonymous):

i got out part 4 as 4 seconds

OpenStudy (anonymous):

I mean part A

OpenStudy (anonymous):

b) x (distance) = v-horizontal. t = 20 x 4 = 80 m

OpenStudy (anonymous):

will vertical distance be same value as horizontal distance in this case ?

OpenStudy (anonymous):

c) Vertical velocity of pebble as it hits ground v = sqrt(2gh) v = sqrt(2 x 10 x 80) v = sqrt(1600) v = 40 m/s

OpenStudy (anonymous):

@Acer100 No, not always the same. It just happens to be the only

OpenStudy (anonymous):

Ok. I see.

OpenStudy (anonymous):

d) v-horizontal is constant, so v-horizontal is equal to 20 m/s

OpenStudy (anonymous):

Thanks alot !!

OpenStudy (anonymous):

ur welcome @Acer100 good luck :)